The number of lost bags per flight at Ango Airport follows a Poisson distribution with population mean. What is the probability that at least 2bags will be lost in a randomly selected flight?

1
Expert's answer
2021-11-17T00:22:03-0500

Let "X=" the number of lost bags per flight: "X\sim Po(\lambda)"

Given "\lambda=0.4"


"P(X\geq 2)=1-P(X=0)-P(X=1)"

"=1-\dfrac{e^{-\lambda}\cdot\lambda^0}{0!}-\dfrac{e^{-\lambda}\cdot\lambda^1}{1!}"

"=1-\dfrac{e^{-0.4}\cdot0.4^0}{0!}-\dfrac{e^{-0.4}\cdot0.4^1}{1!}"

"=0.061552"